Need Help on these questions, Please. Thanks!USE THE FOLLOWING INFORMATION TO FILL IN THE BALANCE SHEET BELOW TO ANSWER QUESTIONS 31 through 34 60 Number of shares outstaUSE THE INFORMATION BELOW TO ANSWER THE FOLLOWING 3 QUESTIONS Proxelo, Inc. Balance Sheets For the Years Ending December 31,
Answer #1

Answer:

Total sales $200,000
Credit sales (60% of total sales) 120,000
Average collection period 60days
Formula of Average collection period 360*Account receivable/Credit Sales
Computation of Account receivable

60 =360 * Account receivable/120,000

Account receivable = (120,000*60) / 360

= $20,000

Account receivable $20,000
Total Sales $200,000
Less:Gross profit margin 20% ( 40,000)
Cost of goods sold $160,000
Inventory turnover ratio 4
Formula of Inventory turnover ratio

Cost of goods sold/Inventory

4 = 160,000/Inventory

Inventory = 160,000/4

Inventory = 40,000

Inventory $40,000
Net profit (15 % of sales)

15% of $200,000

= $30,000

Formula of Return on assets

Net profit/Total assets *100

Return on asset

7.5%

7.5 = 30,000/Total assets *100

Total assets = (30,000*100) / 7.5

=$400,000

Total assets $400,000
Account payable 60days
Formula of Account Payable ratio

Account payable days = 360*Account Payable/cost of goods sold

60=(Account Payable*360)/160,000

Account payable = (160,000*60)/360

=$26,667

Dividend payout ratio 80%
Formula of dividend payout ratio

Dividend paid/Net income*100

80 = Dividend paid/30,000 *100

Dividend paid = (30,000*80)/100

= $24,000

Debt ratio 40%
Formula of debt ratio Total liabilities/ Total assets *100
Tota liabilities

Total assets = $400,000(as calculated above)

40 = (Total liabilities/400,000) *100

Total liabilities = 400,000*40/100

$160,000

Total current liabilities

Account Payable + Notes payable + Accruals

$26,667+$10,000+$5,000

=$41,667

Long term debt

Total liabilities - current assets

160,000-41667

= $118,333


Current ratio 1.5 times
Formula of current ratio current assets/ current liabilities
Current assets

current assets/41,667 = 1.5

Current assets = 62,500

Cash

current assets- (Account receivable +Inventory)

62,500-(20,000+40,000)

= $2,500

Net fixed assets

Total assets - current assets

400,000-62,500

$337,500

Retained earning (2019)

Retained earning (2018)+Net income - Dividend paid

$23,700+$30,000-$24,000

$29,700

Capital surplus Balancing figure


Balance sheet for the year ended,2019

Cash $2,500 Notes payable $10,000
Account receivable 20,000 Account payable 26,667
Inventory 40,000 Accruals 5,000
Net fixed assets 337,500 Long term debt 118,333

Common stock $2 at par value

(Number of outstanding shares = 15,000 shares)

30,000
Capital surplus 180,300
Retained earning 29,700
Total assets 400,000 Total liabilities and equity 400,000